If the graph of the function g defined by g (r)-5x2 +3 is translated vertically upward by 7 units, it becomes the graph of a function /. Find the expression for h (r r)
Answer #1

When ever a function translate vertically upwards or down wards it means there is a constant quantity being added in the independent variable i.e. f(x)-f(x) + a

Where \small a is the amount by which the graphs translate, if upwards, its positive, if downwards its negetive

So lets say a graph of ours i.e. g(x) =-5x2 + 3 is translated to become \small h(x)
And the translate is \small 7 units upwards means \small a=7

So here we have-
\small g(x)\overset{a}{\rightarrow}h(x)
=> h(x) = g(x) + 7
h(x)5r2 +3 + 7
\small =>h(x)=-5x^2+10

So the expression for \small h(x) is-
. h(z) =-52.2 + 10
or
·h (z) =-5(z2-2 (can be rewiritten as this also if we take \small -5 common)